OpenStudy (anonymous):

Solve the system y = -x + 6 and y = x - 8 by using graph paper or graphing technology. (2 points)Question 1 options: 1) (1.5 , -6.5) 2) (2 , 2) 3) (10 , 2) 4) (7 , -1)

myininaya (myininaya):

-x+6=x-8 +x +x 6 =2x-8 +8 +8 14=2x -- -- 2 2 x=7

myininaya (myininaya):

y=-7+6=-1

OpenStudy (anonymous):

Draw the graphs. That is what the question wants. See the attachment. We see that the solution is (4)